OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). @jigglypuff314

OpenStudy (anonymous):

-1 -7(-1) - 9?

jigglypuff314 (jigglypuff314):

yes, that's almost the first step! :) it would be (-1)^2 -7(-1) - 9

OpenStudy (anonymous):

jiggly puff is correct

OpenStudy (anonymous):

jiggly puff is correct

OpenStudy (anonymous):

ok:) sorry i was taking care of my dogs

OpenStudy (anonymous):

(-1)^2 -7(-1) - 9 so it would be 1 - 7 - 9 = -15

jigglypuff314 (jigglypuff314):

careful! :) -7(-1) would become +7 so 1 + 7 - 9 = ?

OpenStudy (anonymous):

it would equal 1?

OpenStudy (anonymous):

or -1

jigglypuff314 (jigglypuff314):

-1 would be more correct :P then plug in -1 for x into x - 2 so -1 - 2 = ?

OpenStudy (anonymous):

-3?

jigglypuff314 (jigglypuff314):

Correct! :D Great Job! ^_^

OpenStudy (anonymous):

thank you i have a couple more:/

OpenStudy (anonymous):

Let f(x) = thirty/ 3 x - 2. Find f(-2).

jigglypuff314 (jigglypuff314):

erm... is it (30/3) x -2 or 30 / (3x-2) ?

OpenStudy (anonymous):

-2= 30/3-2 -2 = 30/ 1 -2 = 30 30/-2 = -15?

OpenStudy (anonymous):

it is 30/(3x-2)

jigglypuff314 (jigglypuff314):

hmmm ok then 30/(3(-2) -2) and use order of operations

OpenStudy (anonymous):

30/(3(-2(-2) 30/-12?

jigglypuff314 (jigglypuff314):

30/(3(-2) -2) as in 30/ (-6 - 2) -> 30 / (-8)

OpenStudy (anonymous):

-3?

jigglypuff314 (jigglypuff314):

mmm I don't get a whole number :/

OpenStudy (anonymous):

i get -3.75 but i dont think that its

jigglypuff314 (jigglypuff314):

it's what I got... but if you think it's wrong, then the set up might have not been correct

OpenStudy (the_fizicx99):

f(g(x))) = (x^2 - 7x - 9) - 2 = x^2 - 7x - 11 Plug in -1 in for x and that's f(g(-1))

OpenStudy (the_fizicx99):

Omagawsh /_\ You're on a different problem D:

jigglypuff314 (jigglypuff314):

lol yeah @tHe_FiZiCx99 xD sorry about that

OpenStudy (anonymous):

I got -3.75 with what jiggly wrote

OpenStudy (anonymous):

i did too

OpenStudy (the_fizicx99):

Ye ^ If you're rounding then that would be -3.6 ♦_♦ Depends how you chose to round :3

OpenStudy (anonymous):

i think so i have 2 more questions

OpenStudy (anonymous):

Let f(x) = 7x - 13. Find f-1(x).

jigglypuff314 (jigglypuff314):

try posting it as a new question post so we don't confuse more people :P

OpenStudy (anonymous):

ok i will